### v3

# Context-measure: Contextualizing Metric for Camouflage

Abstract
Camouflage is primarily context-dependent yet current metrics for camouflaged scenarios overlook this critical factor. Instead, these metrics are originally designed for evaluating general or salient objects, with an inherent assumption of uncorrelated spatial context. In this paper, we propose a new contextualized evaluation paradigm, Context-measure, built upon a probabilistic pixel-aware correlation framework. By incorporating spatial dependencies and pixel-wise camouflage quantification, our measure better aligns with human perception. Extensive experiments across three challenging camouflaged object segmentation datasets show that Context-measure delivers more reliability than existing context-independent metrics. Our measure can provide a foundational evaluation benchmark for various computer vision applications involving camouflaged patterns, such as agricultural, industrial, and medical scenarios. Code is available at https://github.com/pursuitxi/Context-measure.
1 Introduction
Evaluation metrics are fundamental to image segmentation community, serving both as a basis for model benchmarking, e.g., MS COCO [25], PASCAL VOC [7] and a catalyst for methodological innovation [35, 28, 32]. As a field evolves, it naturally calls for specialized evaluation metrics that align with its distinctive problem settings and evaluation philosophy. This study takes camouflaged object segmentation (COS) [13, 11, 21, 36, 22] as a representative case.
Unlike generic segmentation, COS requires models to segment main targets that are blended with their context, exhibiting low target-background contrast and high structural homogeneity. However, existing evaluation practice in COS still relies on generic segmentation metrics, originally tailored for salient scenarios [5, 15], such as the $F_{\beta}$ [2, 3], $S_{\alpha}$ [9, 8], and $E_{\phi}$ [12, 10]. These saliency-era metrics assess model performance solely based on the comparison between the model-predicted foreground maps (FM) and the manually annotated ground-truth (GT), without accounting for the visual complexity introduced by the context in which the targets are embedded. This naturally raises a question: “Should saliency and camouflage share the same metrics?” Before presenting the answer, please refer to Fig. 1: two morphologically identical butterflies exhibit completely different visual characteristics – the red butterfly is camouflaged, while the blue butterfly is salient. The primary reason for this discrepancy lies in the different contexts in which they are embedded. Based on this example, we can conclude that camouflage is inherently context-dependent, whereas these saliency-era metrics are context-independent. At this point, the answer to the question becomes clear: camouflage requires context-dependent evaluation metrics, distinct from those designed for saliency.
Figure 1: Camouflage vs. Saliency. The red butterfly rests within a cluster of red maple leaves, blending with its background, while the blue butterfly perches amid yellow maple leaves, standing out clearly. The ground-truth mask is shown in Appendix.
Our observation drives us to propose a new contextualized paradigm, termed Context-measure, for evaluation in camouflaged scenarios. To achieve contextualization, a fundamental prerequisite is that a correlation can be defined between any two pixels. This is because if all pixel pairs were independent, the target and its surroundings would become semantically disconnected, and the very distinction between “camouflage” and “saliency” would no longer hold. However, existing metrics either rely on the pixel-independence assumption (e.g., IoU, $F_{\beta}$, $E_{\phi}$) or consider only limited dependencies between specific pixel pairs (e.g., $F_{\beta}^{\omega}$, $S_{\alpha}$), both of which fail to meet this fundamental requirement.
To this end, we first establish a pixel correlation framework to model a general metric, upon which we further contextualize it to capture the inherent semantic dependencies within the camouflaged scenario. Specifically, the general framework simulates the human perceptual loop, which consists of two complementary components: the forward inference $F(\text{Ground-Truth}|\text{Prediction})$ and the reverse deduction $R(\text{Prediction}|\text{Ground-Truth})$. Both respectively characterize understanding what the prediction conveys about reality and verifying how reality is reflected in the prediction.
Then, we quantify the pixel-wise camouflage degree and incorporate it as a weighting factor, achieving contextualizing under camouflaged scenarios.
To evaluate the cognitive alignment of metric prediction with human perception, we curate a new dataset containing 600 human-ranked predictions. Compared with existing metrics, our measure achieves a 41% relative improvement in consistency with human judgment. Additionally, we conduct extensive experiments on three commonly used COS benchmarks. Results demonstrate that our Context-measure achieves superior results across four meta-measures [27].
In summary, our contributions are three-fold.
  • We identify the fundamental limitations of existing context-independent metrics and establish a general pixel correlation framework, providing a unified foundation for contextual evaluation.
  • We introduce a pixel-level camouflage degree quantification mechanism that enables fine-grained weighting of object regions according to their concealment level.
  • We propose a contextual evaluation paradigm, and further develop Context-measure – the first COS-specific metric, extensively validated on multiple challenging benchmarks to demonstrate its effectiveness.
2 Related Studies
2.1 Revisiting Saliency-era Metrics
Existing metrics fall into three groups by their visual granularity, with their key features summarized in Table 1.
Pixel-aware Metrics. These assess per-pixel accuracy by directly comparing the predicted FM with the GT. Mean absolute error (MAE) measures the average absolute FM-GT difference at the pixel level. Intersection over Union (IoU) quantifies the overlap between FM and GT, reflecting how accurately the model captures object shape and spatial extent. The F-measure ($F_{\beta}$) [2, 3], defined as the harmonic mean of precision and recall, balances these two aspects to evaluate overall performance. When $\beta=1$, it reduces to the F1 score, which is mathematically related to the Jaccard Index (JI) [20]. Notably, they satisfy the relation of $\text{IoU}=\text{JI}=\frac{\text{F1}}{2-\text{F1}}$, thus yielding identical rankings in practice and often being used interchangeably. The multiscale IoU (mIoU) [1] extends IoU by incorporating multiple resolutions, enabling more comprehensive evaluation across scales and improving sensitivity to fine boundary details.
Region-aware Metrics. Unlike pixel-aware metrics, which assume all pixels are equally important, region-aware metrics address two perceptual flaws: E-Flaw & D-Flaw as in Table 1. The weighted F-measure ($F_{\beta}^{\omega}$) [27] mitigates these issues through distance-based weighting and error dependency modeling. The Structure-measure ($S_{\alpha}$) [9, 8] integrates both region- and object-aware components, providing a more holistic evaluation of structural integrity in FMs.
Image-aware Metrics. These utilize global statistics to evaluate the holistic FM-GT alignment. For instance, the enhanced-alignment measure ($E_{\phi}$) [10, 12] combines local pixel-level alignment with global image-level statistics, yielding results that are more consistent with human visual perception of foreground segmentation quality.
Table 1: Comparison of metric properties for foreground evaluation. Asm means whether the prediction and ground-truth are evaluated asymmetrically. E-Flaw (equal-importance flaw [27]) refers to the metric penalizes all erroneous pixels equally, regardless of perceptual significance. D-Flaw (dependence flaw [27]) assumes pixel-wise independence, thus ignoring spatial correlations. Image/Region/Pixel specify the perceptual hierarchy where a metric operates. Cam denotes the camouflaged scenarios for which the metric was originally designed.
Metric Asm E-Flaw D-Flaw Image Region Pixel Cam
MAE
IoU/JI [20]
$F_{\beta}$ [2, 3]
mIoU [1]
$F_{\beta}^{\omega}$ [27]
$S_{\alpha}$ [8, 9]
$E_{\phi}$ [10, 12]
$C_{\beta}^{\omega}$ (Ours)
  • $\star$ indicates that $S_{\alpha}$ addresses E-Flaw & D-Flaw when erroneous pixels are distributed across regions; otherwise, it fails to handle either.
2.2 Metric flaws for Camouflage
Although aforementioned metrics have their own characteristics, they share several latent flaws when applied to COS.
Context Flaw. Current metrics are inherently grounded in a context-independent assumption, neglecting the spatial correlation between a target and its surrounding areas. As shown in Fig. 2, we present two foreground masks (a1&a2) that are perfectly symmetrical and contain an equal number of pixels. From the perspective of visual perception, mask (a1) exhibits a higher camouflage level than mask (a2) in the corresponding visual area. We observe that two widely-used saliency metrics, $E_{\phi}$ and $S_{\alpha}$, assign identical scores for both masks, i.e., $E_{\phi}=0.75$, $S_{\alpha}=0.70$, while our Context-measure distinguishes them by incorporating the contextual dependence, yielding a higher score for the more camouflaged case (a1).
Correlation Flaw. Most metrics neglect pixel-level correlations by assuming independence either globally or within predefined areas. For instance, MAE, IoU/JI, $F_{\beta}$, and $E_{\phi}$ treat all pixels as independent units; $F_{\beta}^{\omega}$ assumes independence among correctly predicted pixels. $S_{\alpha}$ partially models structural correlations across four quadrants divided by the horizontal and vertical axes through the foreground centroid, yet it still assumes pixel-aware independence within each quadrant. Fig. 2 also shows two foreground masks (b1&b2), each with an equal number of foreground pixels per quadrant, similar to the division used by $S_{\alpha}$. Despite their distinct structural layouts, both $E_{\phi}$ and $S_{\alpha}$ equally treat these masks, i.e., $E_{\phi}=0.41$, $S_{\alpha}=0.45$. In contrast, our Context-measure ($C_{\beta}^{\omega}$) assigns a higher value for mask (b1), because it has relatively better structural representation.
Figure 2: Flaws of context-independent metrics in Camouflage. Compared with two widely-used saliency metrics ($E_{\phi}$ & $S_{\alpha}$), our Context-measure ($C_{\beta}^{\omega}$) more effectively distinguishes masks in camouflaged scenarios, aligning more closely with human perception.
3 The Proposed Context-measure
We introduce a general pixel-correlation scheme (§3.1) that models the human perceptual loop (§3.2): forward inference and reverse deduction between the prediction and GT masks, and then contextualize it to camouflaged scenarios (§3.3).
3.1 Definition of Probabilistic Pixel Correlation
Let $X_{H\times W}$ and $Y_{H\times W}$ denote the predicted and GT masks, where $H$ and $W$ are the image height and width, respectively. For a pixel $p$, we define two foreground pixel sets: $X_{f}:=\{\text{pixel}\ p|X(p)>0\}$ and $Y_{f}:=\{\text{pixel}\ p|Y(p)>0\}$, where $X(p)\in[0,1]$ and $Y(p)\in\{0,1\}$ denote the values of pixel pp. To quantify the correlation between any two pixels, we model their dependency using a probabilistic formulation. Specifically, we estimate the covariance matrix $\mathbf{\Sigma}_{Y}$ from $Y_{f}$, capturing the shape of the GT object. We normalize the covariance as $\hat{\mathbf{\Sigma}}_{Y}=\frac{\alpha^{2}}{\operatorname{Tr}(\mathbf{\Sigma}_{Y})}\mathbf{\Sigma}_{Y}$, where $\alpha$ is a scalar hyperparameter and $\operatorname{Tr}(\cdot)$ denotes the trace operator. This ensures that pixel-wise correlations remain comparable across images regardless of their resolution size. We define the correlation between two pixels conditioned on a specific foreground $Y_{f}$, which is written as:
$P_{Y}(m,n):=\frac{\exp{\Big(-\frac{1}{2}(\textbf{n}-\textbf{m})^{\top}\hat{\mathbf{\Sigma}}^{-1}_{Y}(\textbf{n}-\textbf{m})\Big)}}{2\pi\sqrt{|\hat{\mathbf{\Sigma}}_{Y}|}},$ (1)
where m and n are the coordinates of pixels $m$ and $n$, respectively. $\hat{\mathbf{\Sigma}}^{-1}_{Y}$ and $|\hat{\mathbf{\Sigma}}_{Y}|$ denotes the inverse and determinant of $\hat{\mathbf{\Sigma}}_{Y}$. Using $P_{Y}(\cdot,\cdot)$, we can quantitatively describe how strongly two pixels are related, decaying with their spatial distance according to a Gaussian law. Intuitively, the farther apart two pixels are, the weaker their correlation becomes.
[…]
Figure 3: Our general Context-measure framework. We formulate the foreground evaluation problem into a reciprocal perceptual loop, which alternates between two processes: forward inference $F(Y|X)$ and reverse deduction $R(X|Y)$. Intuitively, the former understands what the prediction conveys about reality, and the latter verifies how reality is reflected in the prediction. Further details are provided in §3.
3.2 Reciprocal Perceptual Loop
In evaluating segmentation quality, human observers typically engage in an iterative comparison – shifting attention back and forth between the prediction and GT masks to assess their perceptual consistency. Such comparison forms a perceptual loop, alternating between forward inference (understanding what the prediction conveys about reality) and reverse deduction (verifying how reality is reflected in the prediction). This inspires us to formulate the general form of Context-measure, $C_{\beta}$, which employs harmonic weighting to integrate the forward inference $F(Y|X)$ and reverse deduction $R(X|Y)$ processes. As illustrated in Fig. 3, through the entire reciprocal perceptual loop, $C_{\beta}$ is defined as:
$C_{\beta}:=\frac{(1+\beta^{2})\cdot F_{m}\cdot R_{m}}{\beta^{2}\cdot F_{m}+R_{m}},$ (2)
where $F_{m}:=\frac{\|F(Y|X)\|_{1}}{\|X\|_{1}}$ and $R_{m}:=\frac{\|R(X|Y)\|_{1}}{\|Y\|_{1}}$ are the forward and reverse terms normalized by its foreground areas, respectively. Here, the scale factor $\beta$ controls the relative importance between $F_{m}$ and $R_{m}$. We next detail the implementation of these two perceptual processes.
Forward Inference. We compare $X$ against $Y$ to understand what the prediction conveys about reality. Based on the pixel-wise correlation defined in Eqn. 1, $F(Y|X)$ is expressed as:
$F(Y|p^{i}):=X(p^{i})\cdot\sum_{j;q^{j}\in Y_{f}}P_{Y}(p^{i},q^{j}),$ (3)
where the $i$-th predicted foreground pixel $p^{i}\!\in\!X_{f}$ is weighted by its correlation to the GT. Considering computational costs, we discretize $P_{Y}(\cdot,\cdot)$ and implement it as a convolution kernel $\mathcal{K}$, following the $3\sigma$ principle to cover more than 99% of the probability mass. Under this approximation, the forward inference is efficiently rewritten as $F(Y|X)\approx X\odot(\mathcal{K}\ast Y)$, where $\odot$ is element-wise multiplication and $\ast$ means convolution, respectively.
Reverse Deduction. This process starts from the $Y$ and verifies how faithfully the prediction reconstructs and reflects the underlying reality. Accordingly, $R(X|Y)$ is defined as:
$R(X|q^{j}):=1-\prod_{i;p^{i}\in X_{f}}\Big[1-X(p^{i})\cdot P_{Y}(p^{i},q^{j})\Big],$ (4)
where $q^{j}\in Y_{f}$ denotes the j-th foreground pixel $q^{j}\in Y_{f}$. Similar to forward inference process, $P_{Y}(\cdot,\cdot)$ is efficiently approximated using a kernel $\mathcal{K}$ centered at $q^{j}$, thus having:
$R(X|q^{j})$ $=1-\exp{\Big(\sum_{i;p^{i}\in X_{f}}\ln{\Big[1-X(p^{i})\cdot P_{Y}(p^{i},q^{j})\Big]}\Big)}$
$\approx 1-\exp{\Big(-\sum_{i;p^{i}\in X_{f}}X(p^{i})\cdot P_{Y}(p^{i},q^{j})\Big)}$ (5)
$\approx 1-\exp{\Big(-\mathcal{K}_{j}\ast X\Big)}.$ (6)
Here, the approximation in Eqn. 5 relies on the first-order Taylor expansion $\ln{(1-x)}\approx-x$, which holds for small values of $x$. The product $X(p^{i})\cdot P_{Y}(p^{i},q^{j})$ is observed to be close to zero across most experimental samples, thus satisfying this approximation. Moreover, since the output of Eqn. 6 lies within $0,1-e^{-1}]$, we apply a normalization factor $\frac{e}{e-1}$ to rescale it to range $[0,1]$. To this end, the reverse process from $Y$ to $X$ becomes:
$R(X|Y)\approx\frac{e}{e-1}\cdot Y\odot\Big[\mathbbm{1}-\exp{\Big(-\mathcal{K}\ast X\Big)}\Big],$ (7)
where $\mathbbm{1}$ is an all-one matrix, and $\exp{(\cdot)}$ is element-wise.
3.3 Contextualizing Context-measure for COS
We extend our generic form of Context-measure (see Eqn. 2) to camouflaged scenarios by quantifying the key visual attribute – camouflaged degree $\mathcal{D}$.
Motivation. Existing camouflage quantification methods [24, 6] typically operate at the object level, measuring how well an entire object blends into its surroundings. However, in COS tasks, camouflage is inherently spatially non-uniform, different parts of the same object may exhibit varying camouflage degrees depending on their local color contrast with the background. This motivates us a pixel-aware quantification in the LAB color space, whose perceptual uniformity better aligns with human visual sensitivity. Compared to RGB, the LAB space decouples luminance and chromaticity, allowing us faithfully estimate color difference [31] between object regions and their surroundings. Next, we detail two consecutive stages for our new method.
Figure 4: Visualization of the Camouflage Degree. Warmer colors (toward red) indicate a higher degree of camouflage, and vice versa.
Figure 5: Illustration of Contextual Overpainting. Object regions are dilated to define contextual surroundings, and overlapping patches are extracted from both regions. Each patch is embedded as a feature vector by combining LAB color values with scaled spatial coordinates. For each object vector, use ANN to find the best contextual match, which is projected back to overpaint the object.
Stage I: Contextual Overpainting. We proceed through this stage in four steps. (a) Extraction. Different from Lamdouar et al. [24], we begin by expanding the object region outward by $k$ pixels through morphological dilation, thereby defining the expanded band as its contextual surroundings. From both the object and its surrounding context, we densely extract all overlapping $N\times N$ patches. (b) Embedding. Each patch is then embedded as a feature vector, obtained by concatenating its flattened LAB color values with standardized spatial coordinates, the latter scaled by a factor $\lambda$ to balance spatial and chromatic contributions. (c) Search. For every object feature vector, we perform an approximate nearest neighbor algorithm (ANN) to search the most compatible feature vector within the contextual surroundings. (d) Overpainting. Finally, the matched feature vector is projected back to the image space and used to overpaint the corresponding object region, resulting in a contextually harmonized reconstruction. This stage is illustrated in Fig. 5.
Stage II: Camouflage Mapping. We calculate $\mathcal{D}$ by measuring the color difference between the original and repainted pixels in the LAB color space. The differences are then fitted with a nonlinear mapping to obtain the camouflage degree:
$\mathcal{D}:=\frac{\exp{\Big(\gamma\cdot(\mathbbm{1}-\frac{\Delta(R_{lab},\ I_{lab})}{100})\Big)}-\mathbbm{1}}{\exp{(\gamma)}-1},$ (8)
where $\gamma$ modulates the nonlinearity. Here, $\Delta$ denotes the CIEDE2000 color difference [31] between the repainted object $R_{lab}$ and the original object $I_{lab}$ in the LAB space. The visualized heatmaps of camouflage degree $\mathcal{D}$ are presented in Fig. 4, where higher values correspond to stronger camouflage, and lower values denote weaker camouflage. We further contextualize metric for camouflage by weighting $R(X|Y)$ (Eqn. 7) with camouflage degree $\mathcal{D}$. Accordingly, the reverse process can be reformulated as:
$R_{w}:=\frac{\|R(X|Y)\odot(Y+\mathcal{D})\|_{1}}{\|Y+\mathcal{D}\|_{1}}.$ (9)
We finally derive the camouflaged form of Context-measure:
$C_{\beta}^{\omega}=\frac{(1+\beta^{2})\cdot F_{m}\cdot R_{w}}{\beta^{2}\cdot F_{m}+R_{w}}.$ (10)
4 Experiments & Analyses
Figure 6: Experimental illustration of meta-study. Meta-Measure 2 (§4.2): an ideal evaluation metric should assign a higher score to FM when the correct GT is used as the reference, compared to when a randomly swapped one is employed (see (a)); Meta-Measure 3 (§4.3): an ideal evaluation metric should assign a higher score to the original FM than to the one with slight random noise added (see (b)); Meta-Measure 4 (§4.4): an ideal evaluation metric for camouflage should maintain stability against slight boundary variations (see (c) & (d)).
In this section, we employ four meta-measures to evaluate the effectiveness of the Context-measure in comparison to existing metrics. An illustrative example is shown in Fig. 6.
[…]
Experimental Setup. All competing metrics were evaluated using their default configurations. In the following experiments, we analyses two variants of our measure. The generic variant, $C_{\beta}$ as defined in Eqn. 2, was implemented with $\alpha=6$ and $\beta=1.0$. The camouflaged variants, $C_{\beta}^{\omega}$ as in Eqn. 10, applied with a contextual band width $k=20$, patch size $N=7$ with an overlap of 3, and hyper-parameters $\alpha=6$, $\beta^{2}=1.2$, $\gamma=8$, and $\lambda=20$.
4.1 Meta-Measure 1: Human Ranking
We argue that humans are the optimal judges for evaluating the quality of segmentation results, and an ideal evaluation metric should exhibit strong consistency with human judgments. To assess this property, we design the first study (abbr. MM#1) to measure the consistency between the metric-induced rankings and human perception outcomes.
$\bullet$ Ranking Data Curation. To the best of our knowledge, there is currently no dataset in the camouflage domain that provides human-annotated rankings for model predictions. To construct such a novel human-ranked camouflage dataset, we proceed in three steps. (a) Image Selection. We randomly sample images from the COD10K test set and segment them using the above seven models. For each sampled image, we select three segmentation results of varying quality from the corresponding model outputs. (b) FM Composition. Each predicted segmentation map is overlaid on the corresponding original image using the alpha channel, with the original image and GT mask provided as references. (c) User Study. We recruited 15 human participants with basic knowledge of visual perception. Each participant was asked to perform pairwise comparisons between samples and cast their votes accordingly. The collective voting results were then validated through discussion to reach a final consensus ranking. In total, from more than 1,000 raw samples, we curated 200 groups of data, each containing three predicted foreground maps, one original image, and one GT mask, forming the final dataset, denoted as CamoHR. A visualization of the dataset construction pipeline is provided in Fig. 7.
Figure 7: Dataset construction pipeline. (a) Image Selection. Random samples from the COD10K test set with three segmentation results of different quality selected from seven models. (b) FM Composition. Each predicted segmentation result is overlaid on the original image via alpha blending, with the image and GT mask as references. (c) User Study. 15 human participants with basic knowledge of visual perception were recruited to perform pairwise comparisons. Only results unanimously agreed upon by all participants are retained as the consensus ranking.
$\bullet$ Evaluation Protocol. We compute the Spearman rank correlation coefficient $\rho\in[-1,1]$ [4] to quantify the consistency. For easier interpretation, we follow [27] in defining $\theta=1-\rho\in[0,2]$, where lower values indicates better ranking consistency. In addition, to facilitate comparative evaluation, this experiment is conducted on two datasets: CamoHR and FMDatabase [10] – the latter being a binary foreground map dataset in salient scenarios that includes human-annotated rankings.
Table 2: Quantitative analysis of evaluation measures across 4 meta-measures. Only IoU results are reported, as IoU, F1, and JI are equivalent according to §2. Additionally, since the FMDatabase does not provide original images, $C_{\beta}^{\omega}$ cannot be computed on this dataset. The best and second-best results are highlighted in red and blue, respectively. $M$: Meta-Measure; DA#1: COD10K [13]; DA#2: NC4K [26]; DA#3: Trans10K [34], $\text{Ours}^{1}$: $C_{\beta}$; $\text{Ours}^{2}$: $C_{\beta}^{\omega}$; 0.01%  : $\leq$0.01%.
MM#1 MM#2 MM#3 MM#4: Erode MM#4: Dilate
Metric CamoHR FMDatabase DA#1 DA#2 DA#3 DA#1 DA#2 DA#3 DA#1 DA#2 DA#3 DA#1 DA#2 DA#3
IoU 7.93% 12.16% 0.05% 0.06% 0.08% 0.55% 0.39% 3.00% 2.76% 1.99% 0.48% 1.67% 1.32% 0.46%
$F_{\beta}$ 17.43% 19.46% 0.34% 0.32% 0.14% 1.48% 2.00% 7.65% 3.28% 2.54% 0.46% 2.31% 1.46% 0.40%
$F_{\beta}^{\omega}$ 5.50% 14.59% 0.09% 0.13% 0.05% 0.01% 0.01% 0.01% 2.12% 1.49% 0.44% 1.13% 0.83% 0.24%
$S_{\alpha}$ 8.25% 13.78% 0.09% 0.06% 0.54% 10.47% 8.43% 0.53% 1.34% 0.97% 0.31% 0.85% 0.77% 0.20%
$E_{\phi}$ 18.00% 12.16% 3.46% 2.37% 1.62% 13.73% 15.64% 5.62% 1.79% 1.16% 0.38% 1.16% 0.93% 0.18%
$\text{Ours}^{1}$ 4.00% 11.35% 0.01% 0.06% 0.03% 0.01% 0.01% 0.01% 1.33% 0.88% 0.30% 0.84% 0.61% 0.15%
$\text{Ours}^{2}$ 3.25% 0.02% 0.03% 0.03% 0.01% 0.01% 0.01% 1.21% 0.80% 0.29% 0.80% 0.64% 0.14%
$\bullet$ Result Analysis. As shown in Table 2, column 3, we reproduce the experiment on the FMDatabase dataset. The results show that existing metrics generally perform well in salient scenarios, exhibiting high consistency with human judgments. However, when applied to camouflage scenes, they exhibit varying degrees of performance fluctuation. Taking $E_{\phi}$ as an example, although it integrates both global statistics and local pixel matching, its global modeling is based merely on the mean energy and lacks the encoding of spatial positional information. In such complex scenes as camouflage, predicted FM can often be globally consistent but locally mismatched. This discrepancy allows background regions to receive undesirably high scores, which leads to evaluation outcomes misaligned with human perception. Notably, our measure, $C_{\beta}$, not only surpasses existing metrics in salient scenarios, but also consistently outperforms them in challenging camouflaged scenarios. Moreover, when contextualizing for camouflage, the $C_{\beta}^{\omega}$ demonstrates superior performance compared to the general $C_{\beta}$, further validating its effectiveness in perceptually aligned assessment.
4.2 Meta-Measure 2: Ground-Truth Switch
In real-world applications, mismatches between predicted FMs and non-corresponding GT masks may occur due to oversights in data processing or evaluation procedures, as shown in Fig. 6 (a). A reliable evaluation metric should be sensitive enough to reflect such semantic mismatches. In other words, even if a predicted foreground map is of high visual quality, its score should drop significantly if the associated GT mask is semantically incorrect. To evaluate this property, we introduce the second meta-measure study (MM#2), which compares the sensitivity of different metrics when confronted with entirely mismatched GT masks.
$\bullet$ Evaluation Protocol. To guarantee the quality of the predicted FMs, we first select samples with an F1 score $\geq$ 0.6 across all experimental data, and then proceed as follows in three steps. (a) Pseudo-GT Sequence. We need to construct a completely deranged sequence of GT masks, where none of the masks corresponds to its original predicted FM. Specifically, a loop is constructed to randomly assign a GT mask to each predicted FM. For each iteration, if the assigned GT does not correspond to the predicted FM (referred to as a pseudo-GT), the loop proceeds; otherwise, the GT is reassigned until a complete mismatch configuration is achieved. (b) Resize Dimension. Then all the pseudo-GTs are resized to match the dimensions of predicted FM. (c) Record Errors. Each predicted FM is evaluated against its correct GT and pseudo-GT, with the error rate of a given metric defined as the proportion of cases where using the pseudo-GT outperforms using the correct GT.
$\bullet$ Result Analysis. As shown in Table 2, columns 4,5,6, most existing metrics exhibit good error detection capabilities for GT semantic mismatches. Similar to the analysis in MM#14.1), $E_{\phi}$ is susceptible to the influence of high similarity background regions, which compromises its ability to detect semantic-level discrepancies.
4.3 Meta-Measure 3: Noise Sensitivity
When slight random noise is introduced into a predicted FM, it may not significantly alter the structural content but can still degrade its visual quality, as shown in Fig. 6 (b). An ideal evaluation measure should be sensitive to such subtle changes and accurately reflect their impact on perceived quality. To assess this property, we design meta-measure study (MM#3) to evaluate the perceptual sensitivity of different metrics to fine-grained variations in the FMs.
$\bullet$ Evaluation Protocol. (a) Image Selection. We again select high-quality predicted FMs (F1 score $\geq$ 0.6) from all experimental data. (b) Noise Addition. We randomly sample points accounting for 1% of the total image pixels within the intersection of the predicted FM and the GT background regions, where Gaussian noise $\epsilon\sim N(0,0.2^{2})$ is added (negative noise values will be truncated). (c) Recode Errors. We evaluate both the noisy and original predictions against the GT, defining the error rate as the proportion of cases where the noisy version outperforms the original, indicating the metric’s sensitivity to fine-grained disturbances.
$\bullet$ Result Analysis. As shown in Table 2, columns 7,8,9, $\text{IoU},F_{\beta},F_{\beta}^{\omega}$ all achieve excellent performance under this meta-study, which is expected given their pixel-level sensitivity, enabling them to capture such subtle errors. Theoretically, $\text{IoU},F_{\beta}$ should perform as impressive as $F_{\beta}^{\omega}$. However, this is not observed in practice. The key reason is that $F_{\beta}^{\omega}$ supports non-binary inputs, whereas $\text{IoU},F_{\beta}$ rely on adaptive thresholding for binarization. When the noise is extremely slight, adaptive thresholding tends to classify such noise as background, yielding identical scores before and after perturbation and thus failing to detect the change. Different from them, $S_{\alpha}$ exhibits significantly poorer performance on the experimental data derived from COD10K and NC4K. This is because that $S_{\alpha}$ divides the image into regions and compares statistical information within each region. For predictions of moderate quality (i.e., from COD10K or NC4K), adding noise may shift regional statistics closer to the GT, occasionally resulting in higher scores. For high-quality predictions (i.e., from Trans10K), even minor noise can disrupt regional structures, allowing $S_{\alpha}$ to demonstrate a level of sensitivity to perturbations that may surpass that of $\text{IoU},F_{\beta}$. Notably, empowered by the pixel correlation framework that models relationships among arbitrary pixels, our measure exhibits exceptional sensitivity to even the slightest noise perturbations, accounting for its superior performance.
4.4 Meta-Measure 4: Annotation Boundary
Figure 8: Challenging case of our camouflage quantification.
For highly camouflaged objects, accurately delineating their boundaries poses a significant challenge. Due to inherent uncertainty at the boundaries, manually annotated GT inevitably involve subjective judgments, which may result in misalignment with the actual object boundaries. We argue that an ideal evaluation metric should be robust to such boundary deviations within a reasonable range, maintaining stability in its evaluation results. To this end, we propose the final meta-measure (MM#4), which aims to assess the stability of each metric under conditions of boundary ambiguity.
$\bullet$ Evaluation Protocol. To simulate the variability between annotations, we applied gentle morphological operations - dilation and erosion - to the GT mask, introducing subtle, non-destructive boundary perturbations. Specifically, we proceed in two steps. (a) Boundary Perturbation. A $3\times 3$ kernel is applied to perform dilation and erosion on the GT masks of all experimental data, respectively. (b) Recode Variation. The variations in each evaluation measure before and after the boundary perturbation are recorded and used as the results of this meta-study.
$\bullet$ Result Analysis. As shown in Table 2, columns 10–15, since only mild and subtle boundary perturbations are introduced, all metrics demonstrate remarkable stability under this setting, with their average variations consistently maintained around the 0.01 level. Additionally, the results also show that the average variation after dilation tends to be smaller than that after erosion. Despite this, our generic form $C_{\beta}$ already outperforms existing metrics, while the camouflage-aware form $C_{\beta}^{\omega}$ performs even better.
Figure 9: Contextualizing for more scenarios.
5 Future Prospects
Although Context-measure demonstrates strong performance in camouflage, there remains a large room to explore.
$\bullet$ Camouflage quantification under color distortion. As presented in Fig. 8, due to applying the artistic stylization (e.g., monochrome filters, strong tone mapping or heavy compression), the scene’s underlying color-space assumption breaks down, resulting in quantification outcomes that do not perfectly align with human perception.
$\bullet$ Contextualizing for more scenarios. Beyond camouflage, our measure can be applied to broader domains. As shown in Fig. 9, without modifying any parameters, we observed that the heatmap generated by our quantification method performs well across various real scenarios [33, 14, 23]. Therefore, we believe our measure is adaptable to diverse applications by tailoring the spatial quantification based on the unique characteristics of task-specific data.
6 Conclusion
In this paper, we analyze existing evaluation metrics in COS and reveal a pervasive issue – the neglect of context. To resolve this, we propose Context-measure, a novel evaluation paradigm that contextualizes evaluation through a probabilistic pixel correlation framework capable of modeling correlation between any pixel pair. Building on the correlation framework, we design an algorithm to quantify pixel-level camouflage degree, enabling practical contextual evaluation. In addition, we construct the CamoHR, a human-ranked camouflage dataset, to assess consistency between metrics and human perception. Extensive meta-measure experiments show that Context-measure outperforms existing metrics, demonstrating superior effectiveness and robustness.
In summary, an ideal evaluation metric should not only serve as a tool for assessing model performance but also act as a driving force for advancing and innovating the field. We hope that our Context-measure can inspire new perspectives and foster progress in camouflage-oriented research.
Supplementary Material
Figure 1: Camouflage vs. Saliency. The red butterfly rests within a cluster of red maple leaves, blending with its background, while the blue butterfly perches amid yellow maple leaves, standing out clearly.
Appendix A Revisiting Saliency-era Metrics
Let the ground-truth mask be denoted as matrix $Y$ and the predicted foreground mask as matrix $X$, where $X,Y\in[0,1]^{m\times n}$. Both $X$ and $Y$ are flattened into $X^{\prime}$ and $Y^{\prime}$, respectively.
$\bullet$ Mean Absolute Error ($M$). It quantifies the average pixel-wise discrepancy between the predicted foreground mask and the ground-truth mask. $M$ is defined as:
$M:=Mean(Y^{\prime}-X^{\prime}).$ (1)
$\bullet$ Intersection over Union $(\text{IoU})$. It evaluates the overlap between prediction and ground truth by computing the ratio of their intersection to their union, serving as a strict region-level similarity metric. The formulation is as follows:
$\text{IoU}:=\frac{\|X\odot Y\|_{1}}{\|X+Y-X\odot Y\|_{1}},$ (2)
where $\odot$ is element-wise multiplication.
$\bullet$ F-measure ($F_{\beta}$). The F-measure [2, 3] is a standard evaluation metric that balances Precision and Recall when comparing the predicted mask with the ground-truth mask. Its computation depends on four commonly used pixel-level quantities: true positives (TPTP) — foreground pixels correctly predicted as foreground; true negatives (TNTN) — background pixels correctly predicted as background; false positives (FPFP) — background pixels incorrectly predicted as foreground; and false negatives (FNFN) — foreground pixels missed by the prediction. These quantities are computed as:
$TP$ $=\|Y\odot X\|_{1},$
$TN$ $=\|(\mathbbm{1}-Y)\odot(\mathbbm{1}-X)\|_{1},$
$FP$ $=\|(\mathbbm{1}-Y)\odot X\|_{1},$
$FN$ $=\|Y\odot(\mathbbm{1}-X)\|_{1},$
where $\mathbbm{1}$ is an all-one matrix, $\odot$ is element-wise multiplication. Using these, Precision and Recall are computed as:
$\text{Precision}=\frac{TP}{TP+FP},\ \text{Recall}=\frac{TP}{TP+FN}.$
Finally, the F-measure combines Precision and Recall through a $\beta$-weighted harmonic mean:
$F_{\beta}:=(1+\beta^{2})\frac{\text{Precision}\cdot\text{Recall}}{\beta^{2}\cdot\text{Precision}+\text{Recall}}.$ (3)
$\bullet$ Weighted F-measure ($F_{\beta}^{\omega}$). The weighted F-measure, proposed by Margolin et al.  [27], addresses the inherent limitations of the F-measure and provides a more faithful quantification of foreground maps. Let the error vector be defined as $E=|Y^{\prime}-X^{\prime}|$. The weighted error vector is computed as:
$E^{\omega}=\min\{E,E\cdot A\}\odot B,$
where $\odot$ is element-wise multiplication, matrix AA models the spatial dependency between pixels, and vector $B$ encodes the varying perceptual importance across locations. Based on $E^{\omega}$, the weighted true positives, true negatives, false positives, and false negatives are computed as:
$TP^{\omega}$ $=\|Y^{\prime}\odot(\mathbbm{1}-E^{\omega})\|_{1},$
$TN^{\omega}$ $=\|(\mathbbm{1}-Y^{\prime})\odot(\mathbbm{1}-E^{\omega})\|_{1}$
$FP^{\omega}$ $=\|(\mathbbm{1}-Y^{\prime})\odot E^{\omega}\|_{1},$
$FN^{\omega}$ $=\|Y^{\prime}\odot E^{\omega}\|_{1},$
where $\mathbbm{1}$ is an all-one matrix. The weighted Precision and Recall are then given by:
$\text{Precision}^{\omega}=\frac{TP^{\omega}}{TP^{\omega}+FP^{\omega}},\ \text{Recall}^{\omega}=\frac{TP^{\omega}}{TP^{\omega}+FN^{\omega}}.$
Finally, the weighted F-measure is defined as:
$F_{\beta}^{\omega}:=(1+\beta^{2})\frac{\text{Precision}^{\omega}\cdot\text{Recall}^{\omega}}{\beta^{2}\cdot\text{Precision}^{\omega}+\text{Recall}^{\omega}}.$ (4)
$\bullet$ Structure measure ($S_{\alpha}$). The $S_{\alpha}$ [8] evaluates the structural consistency between the predicted mask and the ground-truth mask by jointly accounting for object-level integrity and region-level spatial layout. To compute the region-aware component, the image is partitioned into $K$ subregions, and a structural similarity score is computed in each region following an SSIM-inspired formulation. Let $\bar{X}$ and $\bar{Y}$ denote the mean values of $X$ and $Y$, and let $\sigma_{X}$, $\sigma_{Y}$, and $\sigma_{XY}$ be the standard deviations and covariance of $X^{\prime}$ and $Y^{\prime}$. The structural similarity for a given region is defined as:
$ssim(X,Y)=\frac{2\bar{X}\bar{Y}}{\bar{X}^{2}+\bar{Y}^{2}}\cdot\frac{2\sigma_{X}\cdot\sigma_{Y}}{\sigma_{X}^{2}+\sigma_{Y}^{2}}\cdot\frac{\sigma_{XY}}{\sigma_{X}\cdot\sigma_{Y}}.$
The region-aware similarity is then obtained by a weighted aggregation across regions:
$S_{r}=\sum_{k=1}^{K}w_{k}\cdot ssim(X_{k},Y_{k}),$
where $w_{k}$ assigns a region-dependent weight proportional to the amount of ground-truth foreground contained in region $k$. In addition, the object-aware component captures global structural fidelity by separately measuring the consistency of foreground and background regions in the prediction and ground truth. The foreground alignment score ($O_{FG}$) is computed as:
$O_{FG}=\frac{2\cdot Mean(X^{\prime}\odot Y^{\prime})}{(Mean(X^{\prime}\odot Y^{\prime}))^{2}+1+2\lambda\cdot\sigma_{X^{\prime}\odot Y^{\prime}}}.$
The background alignment ($O_{BG}$) is obtained in an analogous manner by substituting $X^{\prime}$ and $Y^{\prime}$ with their complementary forms $(1-X^{\prime})$ and $(1-Y^{\prime})$. The overall object-aware similarity is given by:
$S_{o}=\bar{Y}\cdot O_{FG}+(1-\bar{Y})\cdot O_{BG}.$
Finally, the Structure measure combines the object-aware and region-aware terms as:
$S_{\alpha}:=\alpha\cdot S_{o}+(1-\alpha)\cdot S_{r},$ (5)
where $\alpha\in[0,1]$ controls the relative contributions.
$\bullet$ Enhanced-alignment measure ($E_{\phi}$). The E-measure [10] is a widely used saliency-evaluation metric designed to jointly capture both pixel-level matching and image-level statistics. As a first step, it constructs the bias matrices $\phi_{X}$ and $\phi_{Y}$ for $X$ and $Y$, by subtracting their mean values to remove global intensity bias:
$\phi_{X}=X-Mean(X)\cdot\mathbbm{1},\phi_{Y}=Y-Mean(Y)\cdot\mathbbm{1}.$
Based on these bias matrices, an alignment matrix $\xi$ is computed to measure the per-pixel agreement between the two normalized representations:
$\xi=\frac{2\cdot\phi_{X}\odot\phi_{Y}}{\phi_{X}\odot\phi_{X}+\phi_{Y}\odot\phi_{Y}}.$
Finally, the E-measure aggregates the alignment over the entire image by applying a quadratic enhancement function $(1+\xi)^{2}$ and averaging it across all pixels:
$E_{\phi}:=\frac{1}{4\times m\times n}\|(1+\xi)^{2}\|_{1}.$ (6)